# Gaitsgory’s central functor and the Arkhipov–Bezrukavnikov equivalence in mixed characteristic

### v3

@@ 8. Regular quotient
If $G$ is a type $A$ group, then ${\mathscr{P}}_{0}^{c}({\operatorname{Hk}}_{{\mathcal{I}}})={\mathscr{P}}_{I}^{0}$, and $H=Z_{\hat{G}}(n_{0})$. We do not need this in the proof of the main theorem and will postpone the discussion of this fact to §10 (cf. Proposition 10.8).
@@
The above proposition is the mixed characteristic analogue of a particular case of [Bez04, Proposition 1, Theorem 3]. We sketch the proof here and refer further details to loc.cit. Note that we can regard the regular representation ${\mathcal{O}}(\hat{G})$ of the dual group as a ring object in $\mathrm{Ind}({\operatorname{Rep}}(\hat{G}))$. Then ${\mathscr{Z}}_{0}({\mathcal{O}}(\hat{G}))$ is a ring object in $\mathrm{Ind}({\mathscr{P}}_{0}^{c}({\operatorname{Hk}}_{{\mathcal{I}}}))$. Zorn’s lemma implies that there exists a maximal left ideal subobject ${\mathscr{J}}\subset{\mathscr{Z}}_{0}({\mathcal{O}}(\hat{G}))$, whose quotient will be denoted by ${\mathcal{O}}(H)$. The centrality of ${\mathfrak{Z}}^{0}$ (cf. Lemma 8.2) implies that ${\mathcal{O}}(H)$ is also a ring object. Thus, we define ${\mathcal{O}}(H)$-$\mathrm{Mod}$ as the category of left ${\mathcal{O}}(H)$-modules in $\mathrm{Ind}({\mathscr{P}}_{0}^{c}({\operatorname{Hk}}_{{\mathcal{I}}}))$. Clearly ${\mathcal{O}}(H)$ is a simple object in the abelian category ${\mathcal{O}}(H)$-$\mathrm{Mod}$. Hence, its endomorphism ring $K:=\mathrm{End}_{{\mathcal{O}}(H)}({\mathcal{O}}(H))$ is a division algebra, and $V\mapsto V\otimes K$ defines an equivalence between the category of right finite $K$-modules and the full subcategory in ${\mathcal{O}}(H)$-$\mathrm{Mod}$ generated by ${\mathcal{O}}(H)$ under finite direct sums and subquotients. Now, we deduce that
@@
because the left hand side is a countable $\bar{\mathbb{Q}}_{\ell}$-vector space and hence it must be algebraic. Now, we construct a monoidal fiber functor to invoke the Tannakian formalism.
@@
If $G$ is of type $A$, then the nilpotent element $n_{0}$ is regular.
@@
Our argument is similar to the one in [AR, Section 6.5.8] and uses weight theory. Recall that in 4.27, we posited that the mixed sheaves ${\mathscr{Z}}_{\mathrm{mix}}(V)$ ought to be monodromy-pure of weight 00, as this is the case in equicharacteristic due to a theorem of Gabber whose proof was written up by Beilinson–Bernstein [BB93]. By Proposition 4.28, we know that this holds for minuscule representations. Note that the functor $\Pi^{0}\colon{\mathcal{P}}({\operatorname{Hk}}_{\mathcal{I}})\to{\mathcal{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$ in (8.1) admits a mixed variant $\Pi^{0}_{\mathrm{mix}}$, namely the quotient of mixed perverse sheaves by the ones with positive dimensional support. We claim that the images under $\Pi^{0}_{\mathrm{mix}}$ are monodromy-pure of weight 00. In other words, we want to show that the weight filtration obtained on ${\mathscr{Z}}_{0}(V)$ via push-pull coincides with the monodromy filtration induced by $n_{0}$. It suffices to prove this when $G$ is adjoint, and then we can check that both filtrations are monoidal on $V$, so we can propagate the claim starting from the minuscule case.
@@ 9. Proof of the AB equivalence
If $G$ is of type $A$, then the functor ${\mathscr{F}}_{{\mathcal{I}}{\mathcal{W}}}$ is an equivalence.
@@
If $G$ is of type $A$, then the functor $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$ from 6.7 is an equivalence of abelian categories.
@@
The left side identifies with $(V\otimes H^{n}(\hat{{\mathcal{N}}}_{\mathrm{Spr}},{\mathcal{O}}({\lambda})))^{\hat{G}}$. The higher cohomology of ${\mathcal{O}}_{\hat{{\mathcal{N}}}_{\mathrm{Spr}}}({\lambda})$ vanishes, meaning we only need to consider the right side when $n=0$. Since there exists an equivariant embedding ${\mathcal{O}}_{\hat{{\mathcal{N}}}_{\mathrm{Spr}}}({\lambda})\to W\otimes{\mathcal{O}}_{\hat{{\mathcal{N}}}_{\mathrm{Spr}}}$ for a certain $W\in{\operatorname{Rep}}_{\Lambda}\hat{G}$, the claim reduces to the preceding lemma. ∎
@@
Let us compute the dimension of the right side. After convolution on the right with ${\mathscr{I}}_{-\nu}({\Lambda})$, it vanishes if $n\neq 0$ by the tilting property of ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$, see 7.4, and has otherwise dimension equal to that of the weight space $V(-\nu)$. As for the left side, we have already checked its vanishing if $n\neq 0$ and it has otherwise dimension equal to that of the weight space $V(\nu)$, as one checks via the cohomology of ${\mathcal{O}}(\nu)$ on the Springer resolution $\hat{{\mathcal{N}}}_{\mathrm{Spr}}$, compare with [AR, Subsection 6.6.3]. ∎
@@ 10. Exotic $t$-structure on the Springer resolution
The equivalence in Theorem 9.1 allows us to transport the perverse $t$-structure on ${\mathscr{D}}_{\mathrm{cons}}({\operatorname{Hk}}_{{\mathcal{I}}{\mathcal{W}}})$ to a $t$-structure which we call the exotic t-structure on $\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])$, at least when $G$ is of type $A$. The exotic t-structure has been intrinsically studied in [Bez06, MR16]. In this section, we discuss the exotic t-structure obtained via our object ${\mathrm{Fl}}_{\mathcal{I}}$, and explain how it will be used to prove the assertions in Remark 8.4 for type $A$ groups.
@@
If $G$ is of type $A$, there are isomorphisms
@@
Assume $G$ is of type $A$. Then, the following hold:
@@
    $\Delta_{\nu}^{\mathrm{ex}}|_{{\mathcal{O}}_{\mathrm{r}}}\simeq{\mathcal{O}}(\nu^{-})|_{{\mathcal{O}}_{\mathrm{r}}},$ (10.9)
    where ${\mathcal{O}}_{\mathrm{r}}\subset\hat{{\mathcal{N}}}$ is the regular orbit, and $\nu^{+}$ (resp. $\nu^{-}$) is the dominant (resp. anti-dominant) $W_{\mathrm{fin}}$-conjugate of $\nu$.
@@
    Assume $G$ is of type $A$. Then, for any $\nu\in{\mathbb{X}}_{\bullet}$, there is an isomorphism
@@
$L_{\nu}^{\mathrm{ex}}|_{{\mathcal{O}}_{\mathrm{r}}}\simeq\left\{\begin{aligned} &{\mathcal{O}}&\mathrm{if}\,\ell(w_{\nu})=0\\ &0&\mathrm{otherwise}\end{aligned}\right.$ (10.11)
@@
This is [AR, Proposition 7.2.4] and we sketch it here for type $A$ groups. If $\ell(w_{\nu})=0$, then $\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}}\cong\mathrm{IC}_{\nu}^{{\mathcal{I}}{\mathcal{W}}}$, and we conclude the proof by Lemma 10.3. In general, there exists a unique $\mu\in{\mathbb{X}}_{\bullet}$ such that $\ell(w_{\mu})=0$ and ${\mathrm{Fl}}_{{\mathcal{I}},w_{\nu}}$ and ${\mathrm{Fl}}_{{\mathcal{I}},w_{\mu}}$ belong to the same connected component. Then the proof of Lemma 6.5 and Proposition 6.6 imply that $\mathrm{IC}_{w_{\mu}}^{{\mathcal{I}}{\mathcal{W}}}$ is a composition factor of $\nabla_{\mu}^{{\mathcal{I}}{\mathcal{W}}}$. Then Lemma 10.3 yields that $L_{\mu}^{\mathrm{ex}}$ is a composition factor of $\nabla_{\nu}^{\mathrm{ex}}$. Combining Lemma 3.9, Lemma 10.3, Corollary 10.4 and the previous discussion, both $L_{\mu}^{\mathrm{ex}}$ and $\nabla_{\nu}^{\mathrm{ex}}$ restricts to an equivariant $\hat{G}$-line bundle on $\tilde{{\mathcal{O}}}_{\mathrm{r}}$. In particular, as a composition factor of $\nabla_{\nu}^{\mathrm{ex}}$, $L_{\nu}^{\mathrm{ex}}$ restricts to 00. ∎
Denote by $\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])_{\mathrm{nr}}$ the full subcategory of perfect complexes supported on the complement of ${\mathcal{O}}_{\mathrm{r}}$.
@@
This is [AR, Lemma 7.2.7] and we could also prove it for type $A$ groups via Corollary 10.4 and Proposition 10.5. ∎
For the rest of this section, we apply the previous discussion to study the relation between ${\mathscr{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$ and its full subcategory ${\mathscr{P}}^{c}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$, culminating in the proof that they coincide in type $A$ and so do $H\subset Z_{\hat{G}}(n_{0})$ as promised in Remark 8.4. Recall we define the functor
@@
is an equivalence for type $A$ groups $G$.
@@
Assume $G$ is of type $A$. Then, there exists a unique t-exact equivalence of $\infty$-categories
@@
By Corollary 10.2, the restriction functor ${\mathscr{F}}^{\mathrm{as}}:\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])\rightarrow{\mathscr{D}}^{b}({\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}}))$ is $t$-exact with respect to the exotic $t$-structure on the source and the tautological $t$-structure on the target. Then, to prove ${\mathscr{F}}^{\mathrm{r}}$ is $t$-exact, it suffices to show that every simple object in $\mathrm{Coh}([\hat{G}\backslash{\mathcal{O}}_{\mathrm{r}}])$ is the restriction of an exotic coherent sheaf. The verification of the later assertion can be argued entirely on the coherent side as in [AR, Proposition 7.2.6]. ∎
@@ 11. Equivariant coherent sheaves on the nilpotent cone
of the nilpotent cone of the dual group $\hat{G}$, defined over the coefficient field ${\Lambda}=\bar{\mathbb{Q}}_{\ell}$. In this section, we study the category $\mathrm{Coh}([\hat{G}\backslash\hat{{\mathcal{N}}}])$ by establishing a connection with a certain quotient of ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})$ and proving main results of [Bez09] in the mixed-characteristic setting for type $A$ groups.
@@
Assume $G$ is of type $A$. Then, there exists a unique equivalence of $\infty$-categories:
@@
making the following diagram commutes

### v4

@@ 8. Regular quotient
If $G$ has enough minuscules, then ${\mathscr{P}}_{0}^{c}({\operatorname{Hk}}_{{\mathcal{I}}})={\mathscr{P}}_{I}^{0}$, and $H=Z_{\hat{G}}(n_{0})$. We do not need this in the proof of the main theorem and will postpone the discussion of this fact to §10 (cf. Proposition 10.8).
@@
The above proposition is the mixed characteristic analogue of a particular case of [Bez04, Proposition 1, Theorem 3]. We sketch the proof here and refer further details to loc.cit. Note that we can regard the regular representation ${\mathcal{O}}(\hat{G})$ of the dual group as a ring object in $\mathrm{Ind}({\operatorname{Rep}}(\hat{G}))$. Then ${\mathscr{Z}}_{0}({\mathcal{O}}(\hat{G}))$ is a ring object in $\mathrm{Ind}({\mathscr{P}}_{0}^{c}({\operatorname{Hk}}_{{\mathcal{I}}}))$. Zorn’s lemma implies that there exists a maximal left ideal subobject ${\mathscr{J}}\subset{\mathscr{Z}}_{0}({\mathcal{O}}(\hat{G}))$, whose quotient will be denoted by ${\mathcal{O}}(H)$. The centrality of ${\mathscr{Z}}_{0}$ (cf. Lemma 8.2) implies that ${\mathcal{O}}(H)$ is also a ring object. Thus, we define ${\mathcal{O}}(H)$-$\mathrm{Mod}$ as the category of left ${\mathcal{O}}(H)$-modules in $\mathrm{Ind}({\mathscr{P}}_{0}^{c}({\operatorname{Hk}}_{{\mathcal{I}}}))$. Clearly ${\mathcal{O}}(H)$ is a simple object in the abelian category ${\mathcal{O}}(H)$-$\mathrm{Mod}$. Hence, its endomorphism ring $K:=\mathrm{End}_{{\mathcal{O}}(H)}({\mathcal{O}}(H))$ is a division algebra, and $V\mapsto V\otimes K$ defines an equivalence between the category of right finite $K$-modules and the full subcategory in ${\mathcal{O}}(H)$-$\mathrm{Mod}$ generated by ${\mathcal{O}}(H)$ under finite direct sums and subquotients. Now, we deduce that
@@
Here, the first isomorphism is obtained by restriction to $\delta_{0}\subset{\mathcal{O}}(H)$, and it implies that $K$ is of countable $\bar{\mathbb{Q}}_{\ell}$-dimension. For the second isomorphism, we observe that every division algebra of countable dimension over $\bar{\mathbb{Q}}_{\ell}$ is isomorphic to $\bar{\mathbb{Q}}_{\ell}$. Now, we construct a monoidal fiber functor to invoke the Tannakian formalism.
@@
If $G$ has enough minuscules, then the nilpotent element $n_{0}$ is regular.
@@
Our argument is similar to the one in [AR, Section 6.5.8] and uses weight theory. Recall that in 4.27, we posited that the mixed sheaves ${\mathscr{Z}}_{\mathrm{mix}}(V)$ ought to be monodromy-pure of weight 00, as this is the case in equicharacteristic due to a theorem of Gabber whose proof was written up by Beilinson–Bernstein [BB93], compare with [BB93, Theorem 5.1.2]. By Proposition 4.28, we know that this holds for minuscule representations. Note that the functor $\Pi^{0}\colon{\mathcal{P}}({\operatorname{Hk}}_{\mathcal{I}})\to{\mathcal{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$ in (8.1) admits a mixed variant $\Pi^{0}_{\mathrm{mix}}$, namely the quotient of mixed perverse sheaves by the ones with positive dimensional support. We claim that the images under $\Pi^{0}_{\mathrm{mix}}$ are monodromy-pure of weight 00. In other words, we want to show that the weight filtration obtained on ${\mathscr{Z}}_{0}(V)$ via push-pull coincides with the monodromy filtration induced by $n_{0}$. It suffices to prove this when $G$ is adjoint, and then we can check that both filtrations are monoidal on $V$, see [BB93, Lemma 4.1.2], and also respect splittings, so we can propagate the claim starting from the minuscule case by Lemma 7.7.
@@ 9. Proof of the AB equivalence
If $G$ has enough minuscules, then the functor ${\mathscr{F}}_{{\mathcal{I}}{\mathcal{W}}}$ is an equivalence.
@@
If $G$ has enough minuscules, then the functor $\mathrm{av}_{{\mathcal{I}}{\mathcal{W}}}^{\mathrm{as}}$ from 6.7 is an equivalence of abelian categories.
@@
The left side identifies with $(V\otimes H^{n}(\hat{{\mathcal{N}}}_{\mathrm{Spr}},{\mathcal{O}}(\nu)))^{\hat{G}}$. The higher cohomology of ${\mathcal{O}}_{\hat{{\mathcal{N}}}_{\mathrm{Spr}}}(\nu)$ vanishes, meaning we only need to consider the right side when $n=0$. Since there exists an equivariant embedding ${\mathcal{O}}_{\hat{{\mathcal{N}}}_{\mathrm{Spr}}}(\nu)\to W\otimes{\mathcal{O}}_{\hat{{\mathcal{N}}}_{\mathrm{Spr}}}$ for a certain $W\in{\operatorname{Rep}}_{\Lambda}\hat{G}$, the claim reduces to the preceding lemma. ∎
@@
Let us compute the dimension of the right side. After convolution on the right with ${\mathscr{I}}_{-\nu}({\Lambda})$, it vanishes if $n\neq 0$ by the tilting property of ${\mathscr{Z}}_{{\mathcal{I}}{\mathcal{W}}}(V)$, see 7.9, and has otherwise dimension equal to that of the weight space $V(-\nu)$. As for the left side, we have already checked its vanishing if $n\neq 0$ and it has otherwise dimension equal to that of the weight space $V(\nu)$, as one checks via the cohomology of ${\mathcal{O}}(\nu)$ on the Springer resolution $\hat{{\mathcal{N}}}_{\mathrm{Spr}}$, compare with [AR, Subsection 6.6.3]. ∎
@@ 10. Exotic $t$-structure on the Springer resolution
The equivalence in Theorem 9.1 allows us to transport the perverse $t$-structure on ${\mathscr{D}}_{\mathrm{cons}}({\operatorname{Hk}}_{{\mathcal{I}}{\mathcal{W}}})$ to a $t$-structure which we call the exotic t-structure on $\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])$, at least when $G$ has enough minuscules. The exotic t-structure has been intrinsically studied in [Bez06, MR16]. In this section, we discuss the exotic t-structure obtained via our object ${\mathrm{Fl}}_{\mathcal{I}}$, and explain how it will be used to prove the assertions in Remark 8.4 for groups with enough minuscules.
@@
If $G$ has enough minuscules, there are isomorphisms
@@
Assume $G$ has enough minuscules. Then, the following hold:
@@
    $\Delta_{\nu}^{\mathrm{ex}}|_{\tilde{{\mathcal{O}}}_{\mathrm{r}}}\simeq{\mathcal{O}}(\nu^{-})|_{\tilde{{\mathcal{O}}}_{\mathrm{r}}},$ (10.9)
    where $\tilde{{\mathcal{O}}}_{\mathrm{r}}\subset\hat{{\mathcal{N}}}_{\mathrm{Spr}}$ is the isomorphic preimage of the regular orbit ${\mathcal{O}}_{\mathrm{r}}\subset\hat{N}$ under the Springer map, and $\nu^{+}$ (resp. $\nu^{-}$) is the dominant (resp. anti-dominant) $W_{\mathrm{fin}}$-conjugate of $\nu$.
@@
    Assume $G$ has enough minuscules. Then, for any $\nu\in{\mathbb{X}}_{\bullet}$, there is an isomorphism
@@
$L_{\nu}^{\mathrm{ex}}|_{\tilde{{\mathcal{O}}}_{\mathrm{r}}}\simeq\left\{\begin{aligned} &{\mathcal{O}}&\mathrm{if}\,\ell(w_{\nu})=0\\ &0&\mathrm{otherwise}\end{aligned}\right.$ (10.11)
@@
This is [AR, Proposition 7.2.4] and we sketch it here for groups with enough minuscules. If $\ell(w_{\nu})=0$, then $\nabla_{\nu}^{{\mathcal{I}}{\mathcal{W}}}\cong\mathrm{IC}_{\nu}^{{\mathcal{I}}{\mathcal{W}}}$, and we conclude the proof by Lemma 10.3. In general, there exists a unique $\mu\in{\mathbb{X}}_{\bullet}$ such that $\ell(w_{\mu})=0$ and ${\mathrm{Fl}}_{{\mathcal{I}},w_{\nu}}$ and ${\mathrm{Fl}}_{{\mathcal{I}},w_{\mu}}$ belong to the same connected component. Then the proof of Lemma 6.5 and Proposition 6.6 imply that $\mathrm{IC}_{w_{\mu}}^{{\mathcal{I}}{\mathcal{W}}}$ is a composition factor of $\nabla_{\mu}^{{\mathcal{I}}{\mathcal{W}}}$. Then Lemma 10.3 yields that $L_{\mu}^{\mathrm{ex}}$ is a composition factor of $\nabla_{\nu}^{\mathrm{ex}}$. Combining Lemma 3.9, Lemma 10.3, Corollary 10.4 and the previous discussion, both $L_{\mu}^{\mathrm{ex}}$ and $\nabla_{\nu}^{\mathrm{ex}}$ restricts to an equivariant $\hat{G}$-line bundle on $\tilde{{\mathcal{O}}}_{\mathrm{r}}$. In particular, as a composition factor of $\nabla_{\nu}^{\mathrm{ex}}$, $L_{\nu}^{\mathrm{ex}}$ restricts to 00. ∎
Denote by $\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])_{\mathrm{nr}}$ the full subcategory of perfect complexes supported on the complement of $\tilde{{\mathcal{O}}}_{\mathrm{r}}$.
@@
This is [AR, Lemma 7.2.7] and we could also prove it for groups with enough minuscules via Corollary 10.4 and Proposition 10.5. ∎
For the rest of this section, we apply the previous discussion to study the relation between ${\mathscr{P}}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$ and its full subcategory ${\mathscr{P}}^{c}_{0}({\operatorname{Hk}}_{{\mathcal{I}}})$, culminating in the proof that they coincide assuming the existence of enough minuscules and so do $H\subset Z_{\hat{G}}(n_{0})$ as promised in Remark 8.4. Recall we define the functor
@@
is an equivalence for groups $G$ with enough minuscules.
@@
Assume $G$ has enough minuscules. Then, there exists a unique t-exact equivalence of $\infty$-categories
@@
By Corollary 10.2, the restriction functor ${\mathscr{F}}^{\mathrm{as}}:\mathrm{Perf}([\hat{G}\backslash\hat{{\mathcal{N}}}_{\mathrm{Spr}}])\rightarrow{\mathscr{D}}^{b}({\mathscr{P}}_{\mathrm{as}}({\operatorname{Hk}}_{{\mathcal{I}}}))$ is $t$-exact with respect to the exotic $t$-structure on the source and the tautological $t$-structure on the target. Then, to prove ${\mathscr{F}}^{\mathrm{r}}$ is $t$-exact, it suffices to show that every simple object in $\mathrm{Coh}([\hat{G}\backslash{\mathcal{O}}_{\mathrm{r}}])$ is the restriction of an exotic coherent sheaf. The verification of the latter assertion can be argued entirely on the coherent side as in [AR, Proposition 7.2.6]. ∎
@@ 11. Equivariant coherent sheaves on the nilpotent cone
of the nilpotent cone of the dual group $\hat{G}$, defined over the coefficient field ${\Lambda}=\bar{\mathbb{Q}}_{\ell}$. In this section, we study the category $\mathrm{Coh}([\hat{G}\backslash\hat{{\mathcal{N}}}])$ by establishing a connection with a certain quotient of ${\mathscr{P}}({\operatorname{Hk}}_{{\mathcal{I}}})$ and proving main results of [Bez09] in the mixed-characteristic setting for groups with enough minuscules.
@@
Assume $G$ has enough minuscules. Then, there exists a unique equivalence of $\infty$-categories:
@@
making the following diagram commute